OpenStudy (anonymous):

x+10y=2, for y? whould my answer come out to -x+2/10

OpenStudy (amistre64):

group your numerator, but yes

OpenStudy (anonymous):

what i did was (x+10y)+(-x)=2+(-x) x+10y-x=x+2 x-x+10y=x+2 10y=x=2 10y/10=x+2/10 y=-x+2/10 was this correct and do you know of an easier way to solve this problem?

OpenStudy (amistre64):

just get more familiar with the steps and that will allow you to cut most of it out subtract x and divide at all by 10

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

x+10y=2 -x -x ----------- 10y = 2-x /10 /10 ----------- y = (2-x)/10

OpenStudy (anonymous):

did i do something wrong because i have -x+2/10

OpenStudy (amistre64):

does 2-x = -x+2 ??